OpenStudy (lxelle):

Show that the equation tan(45+x) = 2tan (45-x) can be written in the form tan^2x - 6tanx + 1 = 0.

OpenStudy (anonymous):

I assume the best approach to solving this problem is to apply the sum identity of tangent.

OpenStudy (anonymous):

and remember that the tan of 45 is just 1

OpenStudy (anonymous):

I am going to attempt to type out the first step of the problem and you let me know if you see any problems with it. Lets solve this one together.

OpenStudy (anonymous):

\[\frac{ 1+\tan x }{ 1-\tan x } = 2(\frac{ 1-tanx }{ 1+tanx})\]

OpenStudy (anonymous):

Are you confused how I derived the above equation?

OpenStudy (anonymous):

\[1+2tanx+\tan ^{2}x = 2- 4tanx + 2\tan ^{2}x\]

OpenStudy (anonymous):

If you expand a bit future you get the above

OpenStudy (anonymous):

If you move all the parts to one side you will be left with your above answer

OpenStudy (anonymous):

\[-\tan ^{2}x+6tanx-1\ = 0]

OpenStudy (anonymous):

\[-\tan ^{2}+6tanx-1 = 0\]

OpenStudy (anonymous):

Then multiply both sides of the equation by -1

OpenStudy (anonymous):

Any questions?

OpenStudy (lxelle):

wow thanks. haha i actually know where i went wrong alrd. :)